A train leaves a station and travels north at 75 km/h.Two hours later a second train leaves on a parallel track and travels 125 km/h. How far from the station will they meet?
Answer by mananth(16946)

Train I -----------75km/h
Train II-----------125 km/h

Let them meet at distance x km from the starting point
Time taken by train B = x/125
Time taken by Train A = x/75 +2
Time taken by A - time taken by B = 2
x/75 - x/125= 2
LCD = 375
5x-3x= 2*375
2x=750
x= 375 km